# PEA-Diffusion: Parameter-Efficient Adapter with Knowledge Distillation in non-English Text-to-Image Generation

### v1

@@ Abstract
Text-to-image diffusion models are well-known for their ability to generate realistic images based on textual prompts. However, the existing works have predominantly focused on English, lacking support for non-English text-to-image models. The most commonly used translation methods cannot solve the generation problem related to language culture, while training from scratch on a specific language dataset is prohibitively expensive. In this paper, we are inspired to propose a simple plug-and-play language transfer method based on knowledge distillation. All we need to do is train a lightweight MLP-like parameter-efficient adapter (PEA) with only 6M parameters under teacher knowledge distillation along with a small parallel data corpus. We are surprised to find that freezing the parameters of UNet can still achieve remarkable performance on the language-specific prompt evaluation set, demonstrating that PEA can stimulate the potential generation ability of the original UNet. Additionally, it closely approaches the performance of the English text-to-image model on a general prompt evaluation set. Furthermore, our adapter can be used as a plugin to achieve significant results in downstream tasks in cross-lingual text-to-image generation. Code will be available at: https://github.com/OPPO-Mente-Lab/PEA-Diffusion
@@ 1 Introduction
To deal with the problem, recent researchers have developed T2I generative models that accept native languages as inputs without additional translation procedures. Taiyi-Bilingual [37] uses a new text encoder that is initialized with Chinese CLIP checkpoints and then trains only the text encoder part. AltDiffusion [35] proposes a multi-stage training scheme for multilingual T2I generation. This approach has a relatively high training cost, and although it partially solves the problem of generating culture-relevant images for specific languages, the model is prone to catastrophic forgetting. Some scholars train T2I models from scratch with other language data. For instance, ERNIE-ViLG 2.0 [6] is trained with Chinese native data and capable of producing high-quality images consistent with Chinese language culture. This line of work would be prohibitively expensive, exceeding the capabilities of many research labs. Moreover, this approach cannot lead to a community effect, and English models for different downstream tasks or other tasks commonly used in the community often need to be retrained. GlueGen [23] designs a GlueNet module to align the condition input features of the original image generator by introducing a new encoder. Although this approach has a low training cost, its ability to generate culture-related images for multiple languages and images with general prompts is relatively weak.
As such, exploring a parameter-efficient plug-and-play method that adapts the publicly available, English-naive, pretrained T2I models to suit non-English and culture-specific concepts is becoming more and more urgent. Consequently, we concentrate on a critical research question (RQ): Is it possible to devise a low-cost method to acquire a non-English T2I model by adapting the existing pretrained T2I models efficiently?
@@
To this end, we take SD as the baseline, as it has shown surprisingly good results on T2I tasks. To finetune SD with minimal cost, we propose a novel parameter-efficient adapter (PEA) method with knowledge distillation (KD) to make full use of the pretrained SD model, referred to as PEA-Diffusion. To directly take advantage of the pretrained SD model, the pretrained SD is used as a teacher to guide student learning. Specifically, we propose the PEA module to align the representation space of the new condition encoder in the student with that of the teacher’s image generator. PEA merely consists of a light-weight multi-layer perceptron (MLP), greatly simplifying the model construction process and preserving the image generation capabilities of the original SD model. Different from GlueNet [23] which forces text feature alignment of two text encoders at encoder space, we propose KD loss to compel representation alignment at feature maps and logits space from the UNet module between our PEA-Diffusion and pretrained English SD models. We argue that our method boosts the image generator to understand the information from the new text encoder. To avoid catastrophic forgetting and reduce trainable parameters, UNet parameters are kept frozen during the whole training procedure.
@@
Take Chinese SD as an example. We conduct a thorough experimental analysis on KD to help understand the effect of where to perform feature alignment between teacher and student. In addition, we provided intuitive analysis about what parameters should be updated, facilitating discovering the impact of the SD parameters for adapting SD into language-specific T2I models. It should be noted that the method could be applicable for any other non-English T2I generation model.
Our core contributions are summarized as follows:
@@
  • PEA module is plug-and-play; it can be directly applied to various T2I tasks, including LoRA, ControlNet, inpainting, small SD models that have been pruned and models accelerated through strategies such as Latent Consistency sampling [19], etc.
  • PEA-Diffusion can achieve remarkable image generation performance with culture-specific prompts compared with other cross-lingual approaches or multilingual T2I models. It can also retain the general generation ability of the original English diffusion model with slight degradation on evaluation metrics.
2 Related Work
@@ 2.1 Multilingual Text-to-image Generation
Diffusion models have shown remarkable capability of yielding photorealistic and diverse images given textual descriptions, such as GLIDE [21], Stable Diffusion [26], Imagen [29], DALL-E2 [25], and VQ-Diffusion [8]. Despite their powerful generation abilities, they still only understand English inputs, leading to limitations. To support non-English text prompt, Taiyi [38] finetunes the Chinese text encoder with a diffusion model. ERNIE-ViLG [36] trains the T2I models from scratch. AltDiffusion [35] proposes multilingual text encoder to capture multiple language features, which serve as the DM conditions. The model was trained in three stages: text encoder finetuning, K-V matrix finetuning, and full finetuning. GlueGen [23] has designed a GlueNet module that aligns the original conditional input features of the Image Generator by introducing a new encoder. ENSAD [17] proposes a translation-enhanced multilingual T2I generation model by ensembling multilingual T2I and neural machine translation. BDM [18] proposes a backbone-branch network architecture by integrating an English T2I backbone with a Chinese-native semantics injection branch. Despite their success in the non-English I2I generation, these works are either constrained to deal with distinctive language concepts or isolated from English-native T2I communities. At the same time, it is not possible to balance the lower cost of training with the performance of evaluation sets.
2.2 Multilingual CLIP
@@
CLIP uses contrastive language-image pretraining to connect text and image representations, building on massive English text-image sample pairs [24]. Recently, some works have applied teacher learning [2] or metric learning [1] to train multilingual CLIP models, expanding CLIP to other languages. AltCLIP [4] takes this one step further enhancing the language capabilities, which is implemented by teacher learning and contrastive learning by leveraging a pretrained multilingual text encoder, XLMR [5]. mCLIP [3] achieves multilingual CLIP by aligning the CLIP text model and a multilingual text encoder. To effectively transfer CLIP to language-specific scenarios, some researchers train language-specific CLIPs on large-scale native image-text pairs, e.g., Chinese CLIP [33]. In order to promote PEA-Diffusion to effectively understand language-specific concepts, we append the PEA module after the language-specific CLIP or Multilingual CLIP and train it to uncover non-English cultural concepts within the T2I. The goal is to establish a bridge between the new CLIP and T2I core structures, such as UNet using the PEA module.
2.3 Knowledge Distillation.
@@
KD was first proposed by Hinton et at. [9] for model compression. KD transfers the knowledge in the output logits of a teacher model with a large size to a smaller student model without obvious performance degradation. The feature map in the hidden states and attention output space can also be used for KD to help students mimic teacher capability [13, 11]. Masked Generative Distillation [34] introduces a feature masking scheme into KD to improve students’ representation power by forcing them to recover the masked features. Inspired by masked image modeling, the supervised masked KD model is proposed by inserting label information into self-distillation frameworks. However, to the best of our knowledge, KD has not been fully explored for training multilingual T2I generations. In this paper, we apply KD methods to language transfer in T2I and further harness the power of the PEA module through combination to achieve significant improvements.
@@ 3 Method
In this section, we introduce the details of the proposed PEA-Diffusion. Figure 1 illustrates the overall architecture of PEA-Diffusion for cross-lingual transfer in T2I generation. To inherit the knowledge of the pretrained diffusion model, we use the pretrained SD as a teacher to supervise PEA-Diffusion learning through KD. For language-specific T2I generation, we use the corresponding language-specific CLIP as a text encoder to capture text information. The generator in PEA-Diffusion is the same UNet as that used in SD. A light-weight MLP adapter is added on top of the CLIP text encoder in PEA-Diffusion, facilitating knowledge transfer from teacher to student.
@@ 3.1 Preliminary
Specifically, for an input image $x_{0}\in\mathbb{R}^{H\times W\times 3}$, the encoder $\mathcal{E}$ of the auto-encoder transforms it into a latent representation $z_{0}\in\mathbb{R}^{h\times w\times c}$, where $f=H/h=W/w$ is the downsampling factor and $c$ is the latent feature dimension. The diffusion process is then performed on the latent space, where a conditional UNet [28] denoiser $\epsilon_{\theta}$ is employed to predict noise $\epsilon$ with current timestep $t$, noisy latent $z_{t}$ and generation condition $C$. In text-to-image scenarios, the condition $C=\tau_{\theta}(y)$ is produced by encoding the text prompts $y$ with a pretrained CLIP [24] text encoder $\tau_{\theta}$. Therefore, the overall training objective of SD is defined as
@@
3.2 Motivation
A naive method for non-English text driven image generation is to replace CLIP text encoder of SD with language-specified CLIP encoder. The text encoder is one of the key components of T2I models, as generation requires precise and fine-grained text embedding for guidance. Whereas, modifying an existing text model to another language-specific one is challenging. Because existing text encoder and image generator in SD is tightly coupled, simply replacing another condition encoder can not work. The essential reason is the mis-alignment between the new text encoder and the old image generator. Besides, we experimentally find joint finetuning highly relies on high-quality text-image pairs and easily falls short because of catastrophic forgetting when updating well-trained parameters. Based on these observations, we freeze the trained parameters of the pretrained model and design a lightweight adapter with KD to align new text encoder and image generator.
@@ 3.3 Cross-Lingual Transfer
Objective Given sets of parallel corpus $X^{T}=\{x_{i}^{t}\}_{i=1}^{N}$ and $X^{S}=\{x_{i}^{t}\}_{i=1}^{N}$ and ${X^{S}}^{{}^{\prime}}=\{x_{i}^{t}\}_{i=1}^{M}$ and the corresponding images $I=\{img\}_{i=1}^{N}$ with $N$ denoting the total number of paired samples $X^{T}$ represents the English corpus, while $X^{S}$ represents the non-English corpus, which corresponds to the English corpus, and ${X^{S}}^{{}^{\prime}}$ represents the non-English corpus with relevant cultural. The teacher model $F^{t}$ projects text-to-image pairs $(X^{T},I)=\{(x_{i}^{t},img_{i})\}_{i=1}^{N}$ into source middle feature maps $E^{T}=\{e_{i}^{T}\}_{i=1}^{N}$ and logits output $L^{T}=\{l_{i}^{T}\}_{i=1}^{N}$. Likewise, the student model $F^{s}$ maps $(X^{S},I)$ into middle target feature maps $E^{S}=\{e_{i}^{s}\}_{i=1}^{N}$ and target logits output $L^{S}=\{l_{i}^{S}\}_{i=1}^{N}$ Intuitively, there exists severe distribution mismatch between the source and target logits output, as well as between $E^{S}$ and $E^{T}$ due to the different languages. Therefore, our target is to make $E^{S}$ match with $E^{T}$, so as to transfer teacher knowledge to the student, enabling cross-lingual T2I generation. For the student, we use a language-specific CLIP text encoder instead of the original English text encoder to enhance token- and sentence-level lingual-native representations.
Knowledge Distillation for T2I Generation The original KD leverages Kullback-Leibler (KL) divergence to measures the difference between the categorical probability distribution of the teacher’s logits output and the student’s logits output [9]. However, On the task of T2I generation, since the middle feature $E^{T}$ is the latent value of pixels instead of categorical probability distribution, KL divergence is unable to measure the difference between student and teacher. Consequently, we replace KL divergence with the L2-norm distance between the logits output from student and teacher to extend raw KD for T2I generation, whose loss function is formulated as:
$\mathcal{L}_{LKD}=\mathbb{E}\Big{[}\lVert l_{i}^{s}-l_{i}^{t}\lVert_{2}^{2}\Big{]}$ (2)
$\mathcal{L}_{LKD}$ merely enforces student distribution to match teacher distribution at logits output layer. To further reduce distribution shift of teacher and student, we boost feature alignment of the middle feature maps between $E^{S}$ and $E^{T}$:
$\mathcal{L}_{FLKD}=\mathbb{E}\Big{[}\sum_{l}\lVert e_{i,l}^{s}-e_{i,l}^{t}\lVert_{2}^{2}\Big{]}$ (3)
where $e_{i,l}^{s}$ and $e_{i,l}^{t}$ represent the feature maps of the $l$-th layer in a predefined set of unet layers from the teacher and the student with the $i$-th data samples, respectively. With respective weights $\lambda_{FL}$ and $\lambda_{L}$, the final KD objective is formalized as below:
@@
Adapter Due to the dimension difference between English CLIP and non-English CLIP, we propose a adapter for simple feature transformation to align feature dimension. Moreover, the adapter is trained to learn language-specific information. The adapter is implemented by a MLP with only 6M parameters. To accelerate training and avoid destroying the generalization capability of the pretrained models, we freeze the CLIP encoder and image generator. As a result, the proposed PEA-Diffusion is both parameter- and computation-efficient because just the adapter is trained.
3.4 Training Strategy
For our goal of training a multi-language T2I generation model, we only need to train on English data since both the new CLIP text encoder and the teacher model’s CLIP text encoder can accept English inputs. In this case, we can use aligned KD loss $\mathcal{L}_{KD}$. For our goal of training a language-specific T2I generation model, we adopt a hybrid training strategy, for the target language’s parallel corpus, we only need to use KD loss $\mathcal{L}_{KD}$ to align the distribution. On the other hand, we also have language-specific culture-related data ${X^{S}}^{{}^{\prime}}$. Because there is a significant translation error when translating this data into English and the teacher model cannot provide high-quality knowledge guidance for this type of data, we only use the original loss $\mathcal{L}_{SD}$ for this data boosting the culture understanding ability of the proposed method.

### v2

@@ Abstract
Text-to-image diffusion models are well known for their ability to generate realistic images based on textual prompts. However, the existing works have predominantly focused on English, lacking support for non-English text-to-image models. The most commonly used translation methods cannot solve the generation problem related to language culture, while training from scratch on a specific language dataset is prohibitively expensive. In this paper, we are inspired to propose a simple plug-and-play language transfer method based on knowledge distillation. All we need to do is train a lightweight MLP-like parameter-efficient adapter (PEA) with only 6M parameters under teacher knowledge distillation along with a small parallel data corpus. We are surprised to find that freezing the parameters of UNet can still achieve remarkable performance on the language-specific prompt evaluation set, demonstrating that PEA can stimulate the potential generation ability of the original UNet. Additionally, it closely approaches the performance of the English text-to-image model on a general prompt evaluation set. Furthermore, our adapter can be used as a plugin to achieve significant results in downstream tasks in cross-lingual text-to-image generation.
Keywords:
Diffusion models Parameter-efficient Plug-and-play
footnotetext: Corresponding authors.
@@ 1 Introduction
To address this issue, recent researchers have created T2I generative models that process native languages directly, bypassing any need for translation. Taiyi-Bilingual [38] incorporates a novel text encoder initialized with Chinese CLIP checkpoints, focusing only on training the text encoder. AltDiffusion [36] introduces a multi-stage training process for multilingual T2I generation, which is relatively costly. Some experts begin training T2I models from scratch using different language data. For example, ERNIE-ViLG 2.0 [6] is trained with native Chinese content and can generate high-quality images that reflect Chinese cultural elements. However, this approach is very expensive and often beyond the reach of many research facilities. Additionally, this method does not foster a community impact, necessitating the retraining of English models for various downstream or commonly used community tasks. GlueGen [24] introduces a GlueNet module to align the input features of the original image generator by adding a new encoder. Despite its low training cost, its effectiveness in generating culturally relevant images in multiple languages and handling broad prompts is comparatively limited.
@@
As such, exploring a parameter-efficient plug-and-play method that adapts the publicly available, English-naive, pretrained T2I models to suit non-English and culture-specific concepts is becoming more and more urgent. Consequently, we concentrate on a critical research question (RQ): Is it possible to devise a low-cost method to acquire a non-English T2I model by adapting the existing pretrained T2I models efficiently? To this end, we take SD as the baseline, as it has shown surprisingly good results on T2I tasks. To finetune SD with minimal cost, we propose a novel parameter-efficient adapter (PEA) method with knowledge distillation (KD) to make full use of the pretrained SD model, referred to as PEA-Diffusion. To take advantage of the pre-trained SD model directly, the pre-trained SD is used as a teacher to guide student learning. Specifically, we propose the PEA module to align the representation space of the new condition encoder in the student with that of the teacher’s image generator. PEA merely consists of a light-weight multi-layer perception (MLP), greatly simplifying the model construction process and preserving the image generation capabilities of the original SD model. Different from GlueNet [24] which forces text feature alignment of two text encoders at encoder space, we propose KD loss to compel representation alignment at feature maps and logits space from the UNet module between our PEA-Diffusion and pretrained English SD models. We argue that our method boosts the image generator to understand the information from the new text encoder. To avoid catastrophic forgetting and reduce trainable parameters, UNet parameters are kept frozen during the entire training procedure.
@@
Take Chinese SD as an example. We conduct a thorough experimental analysis on KD to help understand the effect of where to perform feature alignment between the teacher and the student. In addition, we provided intuitive analysis about what parameters should be updated, facilitating discovering the impact of the SD parameters for adapting SD into language-specific T2I models. It should be noted that the method could be applicable for any other non-English T2I generation model. Our core contributions are summarized as follows:
@@
  • PEA module is plug-and-play, it can be directly applied to various T2I tasks, including LoRA, ControlNet, Inpainting, compressed SD models that have been pruned and models accelerated through various strategies.
  • PEA-Diffusion can achieve remarkable image generation performance with culture-specific prompts compared with other cross-lingual or multilingual T2I models. It can also retain the general generation ability of the original English diffusion model with slight degradation on evaluation metrics.
2 Related Work.
@@ 2.1 Multilingual Text-to-image Generation
Diffusion models have shown a remarkable capability of yielding photorealistic and diverse images given textual descriptions, such as GLIDE [22], Stable Diffusion [27], Imagen [30], DALL-E2 [26], and VQ-Diffusion [8]. Despite their powerful generation abilities, they still only understand English input, leading to limitations. To support non-English text prompt, Taiyi [39] finetunes the Chinese text encoder with a diffusion model. ERNIE-ViLG [37] trains the T2I models from scratch. AltDiffusion [36] proposes a multilingual text encoder to capture multiple language features, which serve as DM conditions. The model was trained in three stages: text encoder fine-tuning, $W_{K},W_{V}$ fine-tuning, and full fine-tuning. GlueGen [24] has designed a GlueNet module that aligns the original conditional input features of the Image Generator by introducing a new encoder. ENSAD [18] proposes a translation-enhanced multilingual T2I generation model by ensembling multilingual T2I and neural machine translation. BDM [19] proposes a backbone-branch network architecture by integrating an English T2I backbone with a Chinese native semantics injection branch. Despite their success in the non-English I2I generation, these works are either constrained to deal with distinctive language concepts or isolated from English-native T2I communities. At the same time, it is not possible to balance the lower cost of training with the performance of evaluation sets. IAP [13] proposes to treat images as pivots to establish the connection between the embedding space of different text encoders, but it doesn’t consider to enhance cultural-related generation ability and expand to downstream tasks.
2.2 Multilingual CLIP.
@@
CLIP uses contrastive language-image pretraining to connect text and image representations, based on massive English text-image sample pairs [25]. Recently, some works have applied teacher learning [2] or metric learning [1] to train multilingual CLIP models, expanding CLIP to other languages. AltCLIP [4] takes this one step further enhancing the language capabilities, which is implemented by teacher learning and contrastive learning by leveraging a pretrained multilingual text encoder, XLMR [5]. mCLIP [3] achieves multilingual CLIP by aligning the CLIP text model and a multilingual text encoder. To effectively transfer CLIP to language-specific scenarios, some researchers train language-specific CLIPs on large-scale native image-text pairs, e.g., Chinese CLIP [34]. In order to promote PEA-Diffusion to effectively understand language-specific concepts, we append the PEA module after the language-specific CLIP or Multilingual CLIP and train it to uncover non-English cultural concepts within the T2I. The goal is to establish a bridge between the new CLIP and T2I core structures, such as UNet using the PEA module.
2.2.1 Knowledge Distillation.
@@
KD [9] transfers the knowledge in the output logits of a teacher model with a large size to a smaller student model without obvious performance degradation. The feature map in the hidden states and attention output space can also be used for KD to help students mimic teacher capability [14, 11]. Masked Generative Distillation [35] introduces a feature masking scheme into KD to improve students’ representation power by forcing them to recover the masked features. Inspired by masked image modeling, the supervised masked KD model is proposed by inserting label information into self-distillation frameworks. However, to the best of our knowledge, KD has not been fully explored for training multilingual T2I generations. In this paper, we apply KD methods to language transfer in T2I and further harness the power of the PEA module through combination to achieve significant improvements.
@@ 3 Method
A naive approach for generating images from non-English text is to substitute the CLIP text encoder in SD with a language-specific CLIP encoder. The text encoder is a crucial part of T2I models since exact and detailed text embedding is necessary for accurate generation. However, altering an existing text model to a different language-specific one is difficult. This difficulty arises because SD’s text encoder and image generator are closely integrated, making it ineffective to just swap out the condition encoder. The primary issue is the mismatch between the new text encoder and the old image generator. Furthermore, our experiments show that joint fine-tuning depends heavily on high-quality text-image pairs and often fails due to catastrophic forgetting when updating well-trained parameters. In light of these insights, we keep the pretrained model’s parameters fixed and develop a lightweight adapter with KD to synchronize the new text encoder and the image generator.
Figure 1 illustrates the overall architecture of PEA-Diffusion for cross-lingual transfer in T2I generation. To inherit the knowledge of the pretrained diffusion model, we use the pretrained SD as a teacher to supervise PEA-Diffusion learning through KD. For language-specific T2I generation, we use the corresponding language-specific CLIP as a text encoder to capture text information. The generator in PEA-Diffusion is the same UNet as that used in SD. A light-weight MLP adapter is added on top of the CLIP text encoder in PEA-Diffusion, facilitating knowledge transfer from teacher to student.
@@ 3.1 Preliminary
Specifically, for an input image $x_{0}\in\mathbb{R}^{H\times W\times 3}$, the encoder $\mathcal{E}$ of the auto-encoder transforms it into a latent representation $z_{0}\in\mathbb{R}^{h\times w\times c}$, where $f=H/h=W/w$ is the downsampling factor and $c$ is the latent feature dimension. The diffusion process is then performed in the latent space, where a conditional UNet [29] denoiser $\epsilon_{\theta}$ is employed to predict noise $\epsilon$ with current timestep $t$, noisy latent $z_{t}$ and generation condition $C$. In T2I scenarios, the condition $C=\tau_{\theta}(y)$ is produced by encoding the text prompts $y$ with a pretrained CLIP [25] text encoder $\tau_{\theta}$. Therefore, the overall training objective of SD is defined as
@@ 3.2 Cross-Lingual Transfer
Objective. Given sets of parallel corpus $Y^{T}=\{y_{i}^{t}\}_{i=1}^{N}$ and $Y^{S}=\{y_{i}^{t}\}_{i=1}^{N}$ and ${Y^{S}}^{{}^{\prime}}=\{y_{i}^{t}\}_{i=1}^{M}$ and the corresponding images $I=\{img\}_{i=1}^{N}$ with $N$ denoting the total number of paired samples, $Y^{T}$ represents the English corpus that will be input into the teacher model, $Y^{S}$ represents the non-English corpus with culturally irrelevant content, and ${Y^{S}}^{{}^{\prime}}$ represents the non-English corpus with cultural relevance, they will be input into the student model. The teacher unet denoiser $\epsilon^{T}$ projects pairs $(Y^{T},I)$ into source middle feature maps and source logits output. Likewise, the student unet denoiser $\epsilon^{S}$ projects pairs $(Y^{S},I)$ into target middle feature maps and target logits output. Intuitively, there exists severe distribution mismatch between the source and target output, as well as between $\epsilon^{S}$ and $\epsilon^{T}$ due to the different languages. Therefore, our target is to make $\epsilon^{S}$ match with $\epsilon^{T}$, so as to transfer teacher knowledge to the student, enabling cross-lingual T2I generation. For the student, we use a language-specific CLIP text encoder instead of the original English text encoder to enhance token- and sentence-level lingual-native representations. For brevity, we omit the subscripts of $\mathbb{E}_{\mathcal{E}(x_{0}),\tau_{\theta}(y),\epsilon\sim\mathcal{N}(0,1),t}$ and $z_{t},t$ conditions in the following notations.
Knowledge Distillation for T2I generation. The original KD leverages Kullback-Leibler (KL) divergence to measures the difference between the categorical probability distribution of the teacher’s logits output and the student’s logits output [9]. However, in the task of T2I generation, since the middle feature maps are the latent value of pixels instead of categorical probability distribution, KL divergence is unable to measure the difference between student and teacher. Consequently, we replace KL divergence with the L2-norm distance between the logits output from student and teacher to extend raw KD for T2I generation, whose loss function is formulated as:
$\mathcal{L}_{LKD}=\mathbb{E}\Big{[}\lVert\epsilon^{T}(y^{T})-\epsilon^{S}(y^{S})\lVert_{2}^{2}\Big{]}$ (2)
where $y^{T}\in Y^{T}$ and $y^{S}\in Y^{S}$. $\mathcal{L}_{LKD}$ merely enforces student distribution to match teacher distribution at logits output layer. To further reduce distribution shift of teacher and student, we boost feature alignment of the middle feature maps between $\epsilon^{S}$ and $\epsilon^{T}$:
$\mathcal{L}_{FLKD}=\mathbb{E}\Big{[}\sum_{l}\lVert f(g_{l}^{T}(y^{T}))-f(g_{l}^{S}(y^{S}))\lVert_{2}^{2}\Big{]}$ (3)
where $f$ represents the norm operation, $g_{l}^{T}$ and $g_{l}^{S}$ represent the feature maps of the $l$-th layer in a predefined set of unet layers from the teacher and the student, respectively. With respective weights $\lambda_{FL}$ and $\lambda_{L}$, the final KD objective is formalized as below:
@@
Adapter. Due to the dimension difference between English CLIP and non-English CLIP, we propose an adapter for simple feature transformation to align feature dimension. Moreover, the adapter is trained to learn language-specific information. The adapter is implemented by a MLP with only 6M parameters. To accelerate training and avoid destroying the generalization capability of the pretrained models, we freeze the CLIP encoder and image generator. As a result, the proposed PEA-Diffusion is both parameter- and computation-efficient because just the adapter is trained.
3.3 Training Strategy and Final Objective
For our goal of training a multi-language T2I generation model, we only need to train on English data since both the new CLIP text encoder and the teacher model’s CLIP text encoder can accept English inputs. In this case, we can use aligned KD loss $\mathcal{L}_{KD}$. For our goal of training a language-specific T2I generation model, we adopt a hybrid training strategy, for the target language’s parallel corpus, we only need to use KD loss $\mathcal{L}_{KD}$ to align the distribution. On the other hand, we also have language-specific culture-related data ${Y^{S}}^{{}^{\prime}}$. Because there is a significant translation error when translating these data into English and the teacher model cannot provide high-quality knowledge guidance for this type of data, we only use $\mathcal{L}_{\text{SD}}$ (1) for improving culture understanding ability. The final objective is formalized as follows:
=$\mathcal{L}=\begin{cases}\lambda_{FL}\mathbb{E}\Big{[}\lVert\epsilon^{T}(y^{T})-\epsilon^{S}(y)\lVert_{2}^{2}\Big{]}+\lambda_{L}\mathbb{E}\Big{[}\sum_{l}\lVert f(g_{l}^{T}(y^{T}))-f(g_{l}^{S}(y))\lVert_{2}^{2}\Big{]},&y\in Y^{S}\\ \mathbb{E}\Big{[}\lVert\epsilon-\epsilon^{S}(y)\rVert_{2}^{2}\Big{]},&y\in{Y^{S}}^{{}^{\prime}}\\ \end{cases}$